OpenStudy (anonymous):

Please help me with calc 2 material

OpenStudy (anonymous):

How would I go about finding the Taylor series for:\[f(x) = \int\limits_{0}^{x}((e^t -1)/t) dt\]where a=0?

OpenStudy (anonymous):

Daaawg, hit us up with that problem.

OpenStudy (anonymous):

Find derivatives of the integrand

OpenStudy (anonymous):

Taylor series is: f(a) + f'(a) * (x-a) + (f''(a) * (x-a)^2)/2 + (f'''(a) * (x-a)^3)/(3!)...

OpenStudy (anonymous):

I think the derivative of integrand is:\[(te^t-e^t+1)/t^2\]im not sure what to do next

OpenStudy (anonymous):

Do you know where you Taylor Series should be centered at, If not specified ceter it at 1

OpenStudy (anonymous):

Centered? It gave 0.

OpenStudy (anonymous):

I am still not sure what to do next

OpenStudy (anonymous):

Master, isn't it true that the integral of that function is the antideriative of that function? So what do you get if you take the derivative of the antiderivative?

OpenStudy (anonymous):

The original function?

OpenStudy (anonymous):

Awww dang. You right.

OpenStudy (anonymous):

So that's your first derivative. Evaluate at 0 to get f'(a). Multiply that by (x-a) =x. That gives you the second term of the Taylor series.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

My pleasure =)